I am stuck on these questions below, thanks in advance for your help!
Answer #1

1) X = V*t

it is a equation of straight line

answer = D

2) equation of motion for car A

Sa = Ua t + 0.5a t^2

Sb = Ub t

to over take B

Sa - Sb = 300

-7t +.5*2.4*t^2 = 300

t = 18.994 sec

answer = A

3) V^2 -u^2 = 2a h

u = 0

V = sqrt ( 2*g*h)

V = 18.79

answer = B

4) displacement is the shortest distance of two points

there fore

answer is = A

5) distance travelled by air bag

D = 0.5-( 25*9*10^-3)

D = 0.275m

D = 0.5 * a * t* t

a = 6790.123 m/ sec^2

answer = D
Answer #2

1) D

2) A

3) B

4) A

5) D